OpenStudy (anonymous):

Write the equation of a line with slope 1/2 and including point (4, -1).

OpenStudy (hexagon001):

y=mx +c is equation of a line. the slope is m=0.5. y=-1 and x=4 substitute these values in y=mx+c.. then rearrange to tget c. then you you can write y=mx+c with m=1/2 and c=..whatever the value you find do you get it?

OpenStudy (anonymous):

-1 = .5(4) + c ?

OpenStudy (hexagon001):

yes..now rearrange so you get c=....

OpenStudy (anonymous):

c = -3

OpenStudy (hexagon001):

excellent. now substitute m=0.5 and c=-3 into y=mx+c and there you have the equation of the line

OpenStudy (anonymous):

Thank you

OpenStudy (hexagon001):

my pleasure.. so what is the equation of the line?

OpenStudy (anonymous):

y = 1/2x - 3

OpenStudy (hexagon001):

good. now as a bonus to check if your equation is correct you put x=4 in and you should get y=-1...if you don't then your equation of line is wrong

OpenStudy (anonymous):

I did and got -1 as my answer

OpenStudy (hexagon001):

then the equation of the line is correct.